OpenStudy (anonymous):

solve for x: sqrt{5-sqrt{x+3}}=2

myininaya (myininaya):

\[\sqrt{5 -\sqrt{x+3}}=2 \] square both sides \[5-\sqrt{x+3}=2\] Subtract 5 on both sides \[-\sqrt{x+3}=2-5\] \[-\sqrt{x+3}=-3\] now square both sides

myininaya (myininaya):

\[x+3=9\] not subtract 3 on both sides

myininaya (myininaya):

be sure to plug in the result to check

OpenStudy (anonymous):

shouldn't it be a 4 on your second step

OpenStudy (rogue):

You forgot to square the 2 :P

myininaya (myininaya):

yep

myininaya (myininaya):

\[5-\sqrt{x+3}=4\]*

myininaya (myininaya):

\[-\sqrt{x+3}=4-5\]

myininaya (myininaya):

\[- \sqrt{x+3}=-1\] now square both sides

myininaya (myininaya):

\[x+3=1 => x=-2\] not check this to make sure it is the solution thanks guy for making me aware of my mistake :)

myininaya (myininaya):

now*

OpenStudy (anonymous):

ahh thank you! the step by step way helped a lot!

myininaya (myininaya):

lol great! :)